Find four consecutive even integers with a sum of 196.
A rectangle has width 52 units and perimeter 200 units. Find the length.

Answer by CubeyThePenguin(3113)

consecutive even integers: x, (x+2), (x+4), (x+6)

x + (x+2) + (x+4) + (x+6) = 196
4x + 12 = 196
4x = 184
x = 46

2) length = L

2(52) + 2L = 200
2L = 96
L = 48